If the Fermi energy of silver is 5.5 eV, the wave number of the fastest electron at 0 K has the magnitude in (meter)^-1 is?
1
Expert's answer
2020-06-17T09:17:30-0400

Since, Fermi momentum if

"P_{ F}=\hbar k_{\small F}"

Also,

"P_F=\sqrt{2m_eE_{F}}"

Thus, from above two equation we get,

"k_F=\frac{\sqrt{2m_eE_F}}{\hbar}"

Given,

"E_F=5.5eV=8.81197^{-19}J\\"

Thus,

"k_F\approx 2\times 10^9m^{-1}"